# DyCoke[Uncaptioned image]: Dynamic Compression of Tokens for Fast Video Large Language Models

### v1

@@ Abstract
Video large language models (VLLMs) have significantly advanced recently in processing complex video content, yet their inference efficiency remains constrained because of the high computational cost stemming from the thousands of visual tokens generated from the video inputs. We empirically observe that, unlike single image inputs, VLLMs typically attend visual tokens from different frames at different decoding iterations, making a one-shot pruning strategy prone to removing important tokens by mistake. Motivated by this, we present DyCoke, a training-free token compression method to optimize token representation and accelerate VLLMs. DyCoke incorporates a plug-and-play temporal compression module to minimize temporal redundancy by merging redundant tokens across frames, and applies dynamic KV cache reduction to prune spatially redundant tokens selectively. It ensures high-quality inference by dynamically retaining the critical tokens at each decoding step. Extensive experimental results demonstrate that DyCoke can outperform the prior SoTA counterparts, achieving 1.5$\times$ inference speedup, 1.4$\times$ memory reduction against the baseline VLLM, while still improving the performance, with no training. 11footnotetext: Corresponding author: wanghuan@westlake.edu.cn
@@
Figure 1: Left: We introduce DyCoke (dynamic compression of tokens), a training-free token compression method for fast video large language models. The key innovation of DyCoke over its predecessors is to dynamically remove redundant tokens during the decoding stage, squeezing both the temporal (video frames) and spatial redundancy in visual tokens. Right: Efficiency and performance comparison of various training-free token pruning methods on MVBench [19] with LLaVA-OV-7B [15]. DyCoke surpasses the SoTA counterparts (PruMerge [35], FastV [3]), with 1.5$\times$ inference speedup, 1.4$\times$ memory reduction against the baseline, while still improving the performance.
@@ 1 Introduction
Video large language models (VLLMs) have advanced significantly in understanding diverse video contexts, primarily because of their enhanced reasoning ability for complex multimodal information [20, 21, 45, 19, 18, 42, 15, 39, 4]. Most current VLLMs rely on sequential visual representations. When dozens of video frames are fed into the language model along with a language prompt, the video input is converted into tens of thousands of tokens. Because of the quadratic complexity of attention mechanism, the inherent visual redundancy across video frames leads to a dramatic surge of computational complexity, resulting in prohibitive training and inference costs. Previous research [5, 6, 44, 47] has largely focused on developing lighter large language models (LLMs) with fewer parameters; however, this often significantly diminishes the complex reasoning capabilities of these models. Consequently, how to reduce the number of video tokens while maintaining model performance emerges as a new research direction.
@@
Figure 2: Attention score between the predicted token at different decoding iterations (x-axis) and the input video tokens (y-axis) at the decoding stage of LLaVA-OV-7B [15] (attention score averaged over all attention layers). Note, some video tokens (e.g., frame #1) become less important as the decoding proceeds, while others may instead become more important (e.g., frame #16). This observation motivates us to develop DyCoke, a token compression method that can dynamically exploit the token redundancy during decoding.
Based on the above analysis, we reassess the need for a straightforward and effective token compression method tailored to VLLMs, using the significant temporal and spatial redundancy in video information as a foundational consideration. Thus, in this work, we present the first temporal-spatial dynamic token compression method (DyCoke) to optimize the token representation tailored for VLLMs, as illustrated in Figure 1 (left). The first phase involves designing a plug-and-play, lightweight token compression module that addresses temporal redundancy by merging similar tokens across frames. We group consecutive frames by sampling and identify tokens with overlapping information in adjacent or nearby frames for temporal merging. The second phase maintains a parsimonious KV cache established in the first phase by dynamically pruning less important information, reducing the spatial redundancy visual tokens, while retaining pruned tokens for secondary activations needed for auxiliary computations. Specifically, our approach enables the model to dynamically select a distinct set of tokens at each decoding step, which is essential for preserving performance. Building on this, DyCoke maximizes reasoning ability while substantially reducing visual tokens, resulting in a more streamlined and representative visual token set.
Empirically, the proposed DyCoke demonstrates excellent performance in video reasoning tasks, simplifying visual tokens as much as possible while maintaining model performance, particularly with long encoded inputs, and does not require fine-tuning or parameter modifications. In the first stage, after temporal merging, redundant visual tokens can be adaptively reduced by 50% - 60%. In the second stage, each iteration can dynamically further reduce visual tokens by an additional 70% - 90% based on the first stage. On average, each frame of video input retains 15 tokens for attention matrix calculation, greatly accelerating the inference process. As shown in Figure 1 (right), DyCoke achieves a 1.54$\times$ inference speedup on LLaVA-OV-7B [15], with the lowest memory consumption and highest accuracy. Notably, the method is training-free.
@@
  • We propose a plug-and-play temporal token merging block to effectively reduce the number of visual tokens while preserving key video content information by leveraging temporal redundancy between frames and merging similar tokens.
@@
  • Experimental results on several video inference benchmarks show that DyCoke maintains high inference accuracy and speed while compressing visual tokens and enables the processing of longer video sequences within the same computational budget.
2 Related Works
@@ 2.1 Video Large Language Models
With advances in large language models (LLMs) and their strong multi-modal understanding and reasoning capabilities, many studies have attempted to integrate LLMs with video encoders to leverage these powerful capabilities for video tasks [18, 19, 15, 39, 21, 16, 13, 24, 28, 20, 4]. Representative works, such as VideoChat [18] and VideoLLaMA [20], use video converters to encode video features based on image LLMs, enhancing understanding capabilities by training on extensive video datasets. LLaVA-NeXT-Interleave [16] and LLaVA-OneVision [15] focus on achieving excellent performance across single-image, multi-image, and video scenarios. Although the potential of VLLMs for video understanding and reasoning is being realized, the tens of thousands of visual tokens required for long videos significantly increase both inference time and memory demands. While works such as VILA aim to optimize token usage, substantial hardware resources are still needed for model fine-tuning [23, 39, 13, 20]. Therefore, we desire a token compression method specifically for VLLMs that requires no fine-tuning.
@@ 2.2 Efficient Multi-Modal Large Language Models
While multi-modal large language models (MLLMs) have made significant progress [11, 17, 25, 26, 48], their large-scale training and deployment entail substantial computational costs. LLaVA-1.5 [25, 7, 34] addresses this by using 4-bit rather than 8-bit quantization for compression. MobileVLM [5] and MobilevLM-v2 [6] utilize compact architectures optimized for mobile applications. Additionally, TinyGPT-V [44], LLaVA-Phi [49], and Vary-toy [40] aim to match or exceed the performance of large models by using smaller LLM backbones, such as Phi-2 [12]. MoE-LLaVA [22] employs a mixture of experts to address model sparsity and improve both efficiency and performance. TinyLLaVA [47] explores more lightweight MLLM architectures and training optimizations. However, in most cases, reducing the size of the LMM backbone compromises its reasoning capability. Improving LMMs efficiency by compressing the number of visual tokens offers a promising alternative.
In previous studies, token pruning has been widely adopted to mitigate token redundancy in vision transformers (ViTs) and large language models (LLMs). ToMe [1] propose merging similar tokens within ViTs to consolidate redundant information while preserving task-relevant content across various domains like image, video, and audio processing. TESTA [32] achieves up to a 75% reduction in processed tokens through the use of temporal and spatial aggregation modules. And TempMe [36] address temporal redundancy by progressively merging tokens across neighboring clips. FastV [2] enhances attention efficiency in MLLMs by pruning redundant image tokens based on attention scores, without requiring additional training. LLaVA-PruMerge [35] introduces adaptive token reduction by selecting key visual tokens based on attention scores derived from the CLIP visual encoder. Furthermore, Look-m [38] apply token merging strategies in KV cache to decrease computational costs and support extended multimodal contexts. xGen-MM-Vid [33] maps a sequence of tokens across multiple frames into a compact set of visual tokens, enabling fine-tuning with fewer visual tokens. LazyLLM [9] employ attention maps to progressively prune tokens, reducing time-to-first-token (TTFT). In our study, we present a novel, training-free dynamic token compression strategy specifically designed for VLLMs that fully accounts for the temporal characteristics of visual tokens to preserve model performance as effectively as possible.
@@ 3 Proposed Method
Figure 3: Detailed overview of our DyCoke method. DyCoke compresses visual tokens in VLLMs through a two-stage pruning process: visual token temporal merging (TTM) and KV cache dynamic pruning. Token temporal merging (illustrated in the red dashed box on the left) merges similar tokens in video frames at the prefilling stage, tapping into the temporal redundancy of the video input; KV cache dynamic pruning (illustrated in the blue dashed box on the right) further removes less attended visual tokens in the KV cache dynamically at the decoding stage, exploiting the spatial redundancy in visual tokens. Note, DyCoke is a drop-in training-free approach to accelerate VLLMs.
@@ 3.1 Background on Video LLM Inference
(1) Prefilling Stage. In the prefilling stage, for a video with $M_{v}$ frames, the image encoder maps each frame into $K_{v}$ embedding vectors, $\mathbf{z}_{i}\in\mathbb{R}^{K_{v}\times d_{v}}$, where $d_{v}$ is the dimension of each embedding vector. The $M_{v}$ frames thus form a embedding sequence $\mathbf{Z_{v}}=\left[\mathbf{z}_{1},\mathbf{z}_{2},...,\mathbf{z}_{M_{v}}\right]$, which is fed into the projector, which maps visual tokens into the same feature space as text to facilitate information fusion and alignment across modalities. The projector output is processed to generate the set $H_{v^{\prime}}\in\mathbb{R}^{M_{v}N_{v}\times D}$ of visual tokens, where $N_{v}$ is the token length corresponding to one frame of video.
Simultaneously, the model receives a token sequence prompt $T=\{t_{i}\}_{i=1}^{N_{q}}$, where $t_{i}$ represents the $i$-th token of $N_{q}$ tokens in total. We can get the set of text tokens $H_{q}\in\mathbb{R}^{N_{q}\times D}$ where $D$ is the dimension of hidden state. Next, the visual tokens and text tokens are concatenated as LLM input, $H=\text{concat}[H_{v^{\prime}},H_{q}]$.
@@ 3.2 Our Method: DyCoke
(1) Visual Token Temporal Merging. In VLLMs, the video input contains significant temporal redundancy. Some activities often persist across multiple frames with minimal visual change, while backgrounds and stationary objects frequently contain similar information across frames, causing substantial redundancy. Combining these redundant visual tokens at temporal scales can reduce the total token length of the input, which accelerates VLLMs inference and decreases memory consumption. We introduce a plug-and-play token temporal merging (TTM) module as a first-stage solution to filter out consecutive, redundant visual tokens.
@@
$\mathbf{A}^{(L)}=\mathrm{Softmax}\left(\frac{\mathbf{Q}^{(L)}(\mathbf{K}^{(L)})^{\top}}{\sqrt{D}}\right).$ (5)
We then extract the attention scores of visual tokens and predicted tokens to form a subset $\mathbf{A}_{v}^{(L)}$. To obtain the top $p\%$ attention scores in the cross-attention matrix $\mathbf{A}_{v}^{(L)}$, we calculate a threshold $\tau$ and define the set $\mathcal{I}_{p}^{(L)}$ comprising the indices of these top $p\%$ attention scores. Then we prune the visual tokens in the KV cache, retain tokens with the high attention scores, and update the KV cache:
@@
In the next decoding step, the model may need to refocus on tokens that were previously pruned. If they are discarded directly, the model cannot retrieve their KV cache entries. We also consider that tokens of interest tend to remain consistent across successive iterations. To reduce large-scale indexing requirements, we introduce a dynamic pruning cache (DP cache) to store pruned tokens, which can be denoted as
@@
where $\mathcal{J}^{(L)}=\{i\mid i\notin\mathcal{I}p^{(L)}\}$. In the next decoding iteration, the cross-attention matrix is recalculated at layer $L$, dynamically adding tokens from the DP cache with increased attention scores into the calculation and storing them in the KV cache. Simultaneously, tokens whose attention scores have decreased since the previous stage and no longer fall within the top $p\%$ are indexed. These tokens are removed from the KV cache and stored in the DP cache. This process repeats at each decoding stage, with the KV cache and DP cache dynamically updated for optimal token compression.
@@ 4.1 Evaluation Setups and Implementation Details
Benchmarks. We evaluate the performance of VLLMs using established video-to-text benchmarks. ActivityNet-QA [43] contains human-annotated, action-related QA pairs derived from the ActivityNet dataset. Single-word responses are generated for evaluation. We evaluate the accuracy of the model and response quality (scored from 0 to 5) using GPT-4o-mini [29]. We also utilize the PerceptionTest [30] to assess perception capabilities, alongside VideoMME [8] and NeXTQA [41], which encompass various video domains and durations. VideoDetailCaption [27] assesses the model’s detailed understanding of video content, with scoring also conducted via GPT-4o-mini [29]. Additionally, we further evaluate the model on the MVbench [19] dataset, which includes 20 complex tasks requiring comprehensive video understanding beyond single-frame analysis. Each task contains 200 test samples in a multiple-choice VideoQA format. These samples require the model to choose the correct answer from multiple provided options.
Computing Cost Evaluation. We examine the total FLOPs of the prefilling stage and the decoding stage. For a transformer layer with multi-head attention (MHA) and feed-forward network (FFN) modules, let $n$ represent the number of tokens, $d$ be the hidden state size, and $m$ be the intermediate size of FFN. In the prefilling phase, the total FLOPs can be estimated as $4nd^{2}+2n^{2}d+2ndm$. For the decoding phase, considering the significant contribution of the KV cache, the computational consumption for $R$ total iterations (i.e., predicting $R$ tokens) is $R\left(4d^{2}+2dm\right)+2\sum_{i=1}^{R}d\times(n+i)$. We unify $R=100$ for calculation in the experiments. Thus, for an LLM with $T$ total transformer layers, the total FLOPs can be expressed as follows,
$\mathrm{FLOPs}$ $=T(4nd^{2}+2n^{2}d+2ndm)$ (8)
$+TR\left((4d^{2}+2dm)+2\left(dn+\frac{d(R+1)}{2}\right)\right).$
@@
Comparison Methods. We compare our method with two latest training-free visual token compression methods: LLaVA-PruMerge [35] leverages the sparsity of attention scores in CLIP [31] to identify key tokens; FastV [3] adopts the attention score between the predicted token and the visual tokens during the prefilling stage to identify key tokens. Notably, FastV offers two versions, with and without KV cache. In practice, we found that FastV without KV cache causes a sharp increase in inference time for video input, thereby we mainly compare to the version with KV cache. Of note, the above two methods rely on one-shot token pruning, while our approach introduces dynamic visual token pruning for the first time. We use the official codes [^1] https://github.com/pkunlp-icler/FastV, https://github.com/42Shawn/LLaVA-PruMerge of these methods for evaluation under identical hardware conditions.
@@
Implementation Details. We implement the proposed DyCoke on the LLaVA-OneVision-0.5B, LLaVA-OneVision-7B, and LLaVA-OneVision-72B models using NVIDIA RTX 4090 (24GB), A6000 (48GB), and A100 (80GB) GPUs, respectively. To set the pruning ratios for all methods, we use total calculated FLOPs as the metric to ensure fair comparison. The attention computation layer for FastV is set to layer 5. For video input, we follow the official requirements of the LLaVA-OneVision model, with a default of 32 video input frames and $N_{v}=196$, except for experiments with specific instructions. In the comparison experiments, $L$ is set to 3 and $P$ to 0.7, with the first-stage pruning rate $K$ serving as the primary experimental variable. For benchmarks such as PerceptionTest [30], VideoMME [8], NeXTQA, VideoDetailCaption [27], and ActivityNet-QA [43], we use the LMMs-Eval [46, 14] for evaluation, while MVBench [19] is evaluated using the official code.
@@
Method Total Latency $\downarrow$ GPU Mem. $\downarrow$ Accuracy $\uparrow$ Latency per Example $\downarrow$ LLaVA-OV-7B Full Tokens 1:19:27 27G 57.56 1.19s (1.00$\times$) PruMerge 2:02:17 20G 52.56 1.83s (0.65$\times$) FastV 1:03:50 24G 56.07 0.96s (1.24$\times$) Ours ($K=0.5$) 59:02 21G 57.96 0.88s (1.35$\times$) Ours ($K=0.7$) 57:13 19G 57.50 0.85s (1.40$\times$) LLaVA-OV-0.5B Full Tokens 34:05 19G 47.09 0.56s (1.00$\times$) PruMerge 1:35:08 8.5G 42.51 1.42s (0.41$\times$) FastV 32:30 10G 46.14 0.49s (1.14$\times$) Ours ($K=0.5$) 32:17 8.9G 46.96 0.48s (1.16$\times$) Ours ($K=0.7$) 31:45 7.4G 47.07 0.47s (1.19$\times$) (a) The number of video input frames is 16 Method Total Latency $\downarrow$ GPU Mem. $\downarrow$ Accuracy $\uparrow$ Latency per Example $\downarrow$ LLaVA-OV-7B Full Tokens 2:33:30 34G 58.79 2.3s (1$\times$) PruMerge 3:48:59 28G 53.88 3.43s (0.64$\times$) FastV 1:55:20 30G 58.36 1.73s (1.32$\times$) Ours ($K=0.5$) 1:53:17 28G 59.06 1.69s (1.36$\times$) Ours ($K=0.7$) 1:39:49 24G 59.60 1.49s (1.54$\times$) LLaVA-OV-0.5B Full Tokens 1:19:06 21G 48.23 1.18s (1$\times$) PruMerge 3:06:06 13G 42.55 2.79s (0.42$\times$) FastV 1:13:25 15G 47.01 1.10s (1.07$\times$) Ours ($K=0.5$) 1:10:32 12G 48.10 1.06s (1.11$\times$) Ours ($K=0.7$) 1:02:15 10G 47.79 0.93s (1.27$\times$) (b) The number of video input frames is 32

### v3

@@ Abstract
Video large language models (VLLMs) have significantly advanced recently in processing complex video content. Yet, their inference efficiency remains constrained because of the high computational cost stemming from the thousands of visual tokens generated from the video inputs. We empirically observe that, unlike single image inputs, VLLMs typically attend visual tokens from different frames at different decoding iterations. This makes a one-shot pruning strategy prone to removing important tokens by mistake. Motivated by this, we present DyCoke, a training-free token compression method to optimize token representation and accelerate VLLMs. DyCoke incorporates a plug-and-play temporal compression module to minimize temporal redundancy by merging redundant tokens across frames and applying dynamic KV cache reduction to prune spatially redundant tokens selectively. It ensures high-quality inference by dynamically retaining the critical tokens at each decoding step. Extensive experimental results demonstrate that DyCoke can outperform the prior SoTA counterparts, achieving 1.5$\times$ inference speedup, and 1.4$\times$ memory reduction against the baseline VLLM, while still improving the performance, with no training. 11footnotetext: Corresponding author: wanghuan@westlake.edu.cn
@@
Figure 1: Left: We introduce DyCoke (dynamic compression of tokens), a training-free token compression method for fast video large language models. The key innovation of DyCoke over its predecessors is to dynamically remove redundant tokens during the decoding stage, squeezing both the temporal (video frames) and spatial redundancy in visual tokens. Right: Efficiency and performance comparison of various training-free token pruning methods on MVBench [23] with LLaVA-OV-7B [18]. DyCoke surpasses the SoTA counterparts (PruMerge [39], FastV [3]), with 1.5$\times$ inference speedup and a 1.4$\times$ reduction in memory usage relative to the baseline, while simultaneously enhancing performance.
@@ 1 Introduction
Video large language models (VLLMs) have advanced significantly in understanding diverse video contexts, primarily because of their enhanced reasoning ability for complex multimodal information [24, 25, 52, 23, 22, 49, 18, 43, 4]. Most current VLLMs rely on sequential visual representations. When dozens of video frames are fed into the language model along with a language prompt, the video input is converted into tens of thousands of tokens. Because of the quadratic complexity of the attention mechanism, the inherent visual redundancy across video frames leads to a dramatic surge of computational complexity, resulting in prohibitive training and inference costs. Previous research [5, 6, 51, 54] has largely focused on developing lighter large language models (LLMs) with fewer parameters; however, this often significantly diminishes the complex reasoning capabilities of these models. Consequently, how to reduce the number of video tokens while maintaining model performance emerges as a new research direction.
@@
Figure 2: Attention score between the predicted token at different decoding iterations (x-axis) and the input video tokens (y-axis) at the decoding stage of LLaVA-OV-7B [18] (attention score averaged over all attention layers). Note, that some video tokens (e.g., frame #1) become less important as the decoding proceeds, while others may instead become more important (e.g., frame #16). This observation motivates us to develop DyCoke, a training-free plug-and-play token compression method that can dynamically exploit the token redundancy during decoding.
Based on the above analysis, we reassess the need for a straightforward and effective token compression method tailored to VLLMs, using the significant temporal and spatial redundancy in video information as a foundational consideration. Thus, in this work, we present the first temporal-spatial dynamic token compression method (DyCoke) to optimize the token representation tailored for VLLMs, as illustrated in Figure 1 (left). The first phase involves designing a plug-and-play, lightweight token compression module that addresses temporal redundancy by merging similar tokens across frames. We group consecutive frames by sampling and identifying tokens with overlapping information in adjacent or nearby frames for temporal merging. The second phase maintains a parsimonious KV cache established in the first phase by dynamically pruning less important information, reducing the spatial redundancy of visual tokens, while retaining pruned tokens for secondary activations needed for auxiliary computations. Specifically, our approach enables the model to dynamically select a distinct set of tokens at each decoding step, which is essential for preserving performance. Building on this, DyCoke maximizes reasoning ability while substantially reducing visual tokens, resulting in a more streamlined and representative visual token set.
Empirically, the proposed DyCoke demonstrates excellent performance in video reasoning tasks, simplifying visual tokens as much as possible while maintaining model performance, particularly with long encoded inputs, and does not require fine-tuning or parameter modifications. In the first stage, after temporal merging, redundant visual tokens can be adaptively reduced by 50% - 60%. In the second stage, each iteration can dynamically further reduce visual tokens by an additional 70% - 90% based on the first stage. On average, each video input frame retains 15 tokens for attention matrix calculation, greatly accelerating the inference process. As shown in Figure 1 (right), DyCoke achieves a 1.54$\times$ inference speedup on LLaVA-OV-7B [18], with the lowest memory consumption and highest accuracy. Notably, the method is training-free.
@@
  • We propose a plug-and-play temporal token merging block to effectively reduce visual tokens while preserving the key video content information by leveraging temporal redundancy between frames and merging similar tokens.
@@
  • Experimental results on several video inference benchmarks show that DyCoke maintains high inference accuracy and speed while compressing visual tokens and enables processing longer video sequences within the same computational budget.
2 Related Work
@@ 2.1 Video Large Language Models
With advances in large language models (LLMs) and their strong multi-modal understanding and reasoning capabilities, many studies have attempted to integrate LLMs with video encoders to leverage these powerful capabilities for video tasks [22, 23, 18, 43, 25, 19, 16, 28, 32, 24, 4, 21, 48, 45]. Representative works, such as VideoChat [22] and VideoLLaMA [24], use video converters to encode video features based on image LLMs, enhancing understanding capabilities by training on extensive video datasets. LLaVA-NeXT-Interleave [19] and LLaVA-OneVision [18] focus on achieving excellent performance across single-image, multi-image, and video scenarios. Although the potential of VLLMs for video understanding and reasoning is being realized, the tens of thousands of visual tokens required for long videos significantly increase both inference time and memory demands. While works such as VILA aim to optimize token usage, substantial hardware resources are still needed for model fine-tuning [27, 43, 16, 24]. Therefore, we desire a token compression method specifically for VLLMs that requires no fine-tuning.
@@ 2.2 Efficient Multi-Modal Large Language Models
While multi-modal large language models (MLLMs) have made significant progress [12, 20, 29, 30, 55, 47, 15, 15, 47, 10, 13], their large-scale training and deployment entail substantial computational costs. LLaVA-1.5 [29, 7, 38] addresses this by using 4-bit rather than 8-bit quantization for compression. MobileVLM [5] and MobilevLM-v2 [6] utilize compact architectures optimized for mobile applications. Additionally, TinyGPT-V [51], LLaVA-Phi [56], and Vary-toy [44] aim to match or exceed the performance of large models by using smaller LLM backbones, such as Phi-2 [14]. MoE-LLaVA [26] employs a mixture of experts to address model sparsity and improve both efficiency and performance. TinyLLaVA [54] explores more lightweight MLLM architectures and training optimizations. However, many of these works have noted that reducing the size of the LMM model backbone often compromises its reasoning capability. Improving LMM efficiency by compressing the number of visual tokens offers a promising alternative.
In previous studies, token pruning has been widely adopted to mitigate token redundancy in vision transformers (ViTs) and large language models (LLMs) [57, 3]. ToMe [1] proposes merging similar tokens within ViTs to consolidate redundant information while preserving task-relevant content across various domains like image, video, and audio processing. TESTA [36] achieves up to a 75% reduction in processed tokens through the use of temporal and spatial aggregation modules. TempMe [40] addresses temporal redundancy by progressively merging tokens across neighboring clips. FastV [2] enhances attention efficiency in MLLMs by pruning redundant image tokens based on attention scores, without requiring additional training. LLaVA-PruMerge [39] introduces adaptive token reduction by selecting key visual tokens based on attention scores derived from the CLIP visual encoder. Furthermore, Look-m [42] applies token merging strategies in the KV cache to decrease computational costs and support extended multimodal contexts. xGen-MM-Vid [37] maps a sequence of tokens across multiple frames into a compact set of visual tokens, enabling fine-tuning with fewer visual tokens. LazyLLM [9] employs attention maps to progressively prune tokens, reducing time-to-first-token (TTFT). In our study, we present a novel, training-free dynamic token compression strategy specifically designed for VLLMs that fully accounts for the temporal characteristics of visual tokens to preserve model performance as effectively as possible.
@@ 3 Proposed Method
Figure 3: Detailed overview of our DyCoke method. DyCoke compresses visual tokens in VLLMs through a two-stage pruning process: visual token temporal merging (TTM) and KV cache dynamic pruning. Token temporal merging (illustrated in the red dashed box on the left) merges similar tokens in video frames at the prefilling stage, tapping into the temporal redundancy of the video input; KV cache dynamic pruning (illustrated in the blue dashed box on the right) further removes less attended visual tokens in the KV cache dynamically at the decoding stage, exploiting the spatial redundancy in visual tokens. DyCoke is a drop-in training-free approach to accelerate VLLMs.
@@ 3.1 Background on Video LLM Inference
(1) Prefilling Stage. In the prefilling stage, for a video with $M_{v}$ frames, the image encoder maps each frame into $K_{v}$ embedding vectors, $\mathbf{z}_{i}\in\mathbb{R}^{K_{v}\times d_{v}}$, where $d_{v}$ is the dimension of each embedding vector. The $M_{v}$ frames thus form an embedding sequence $\mathbf{Z_{v}}=\left[\mathbf{z}_{1},\mathbf{z}_{2},...,\mathbf{z}_{M_{v}}\right]$, which is fed into the projector, which maps visual tokens into the same feature space as text to facilitate information fusion and alignment across modalities. The projector output is processed to generate the set $H_{v^{\prime}}\in\mathbb{R}^{M_{v}N_{v}\times D}$ of visual tokens, where $N_{v}$ is the token length corresponding to one frame of video.
Simultaneously, the model receives a token sequence prompt $T=\{t_{i}\}_{i=1}^{N_{q}}$, where $t_{i}$ represents the $i$-th token of $N_{q}$ tokens in total. We can get the set of text tokens $H_{q}\in\mathbb{R}^{N_{q}\times D}$ where $D$ is the dimension of the hidden state. Next, the visual tokens and text tokens are concatenated as LLM input, $H=\text{concat}[H_{v^{\prime}},H_{q}]$.
@@ 3.2 Our Method: DyCoke
(1) Visual Token Temporal Merging. In VLLMs, the video input contains significant temporal redundancy. Some activities often persist across multiple frames with minimal visual change, while backgrounds and stationary objects frequently contain similar information across frames, causing substantial redundancy. Combining these redundant visual tokens at temporal scales can reduce the total token length of the input, which accelerates VLLM inference and decreases memory consumption. We introduce a plug-and-play token temporal merging (TTM) module as a first-stage solution to filter out consecutive, redundant visual tokens.
@@
$\mathbf{A}^{(L)}=\mathrm{Softmax}\left(\frac{\mathbf{Q}^{(L)}(\mathbf{K}^{(L)})^{\top}}{\sqrt{D}}\right),$ (5)
where $\mathbf{Q}^{(L)}\in\mathbb{R}^{1\times d}$. We then extract the attention scores of visual tokens and predicted tokens to form a subset $\mathbf{A}_{v}^{(L)}$. To obtain the top $p\%$ attention scores in the cross-attention matrix $\mathbf{A}_{v}^{(L)}$, we calculate a threshold $\tau$ and define the set $\mathcal{I}_{p}^{(L)}$ comprising the indices of these top $p\%$ attention scores. Then we prune the visual tokens in the KV cache, retain tokens with high attention scores, and update the KV cache:
@@
In the next decoding step, the model may need to refocus on the tokens that were previously pruned. If they are discarded directly, the model cannot retrieve their KV cache entries. We also consider that tokens of interest tend to remain consistent across successive iterations. To reduce large-scale indexing requirements, we employ cosine similarity to measure the attention distribution across different decoding iterations. The KV Cache is updated only at iteration $N$, where a low similarity is observed. We also introduce a dynamic pruning cache (DP cache) to store pruned tokens, which can be denoted as
@@
where $\mathcal{J}^{(L)}=\{i\mid i\notin\mathcal{I}p^{(L)}\}$. In the iteration $N$, the cross-attention matrix is recalculated at layer $L$, dynamically adding tokens from the DP cache with increased attention scores into the calculation and storing them in the KV cache. Simultaneously, tokens whose attention scores have decreased since the previous stage and no longer fall within the top $p\%$ are indexed. These tokens are removed from the KV cache and stored in the DP cache. This process repeats at each decoding stage, with the KV cache and DP cache dynamically updated for optimal token compression.
@@ 4.1 Evaluation Setups and Implementation Details
Benchmarks. We evaluate the performance of VLLMs using established video-to-text benchmarks. ActivityNet-QA [50] comprises human-annotated, action-related question-answer pairs derived from the ActivityNet dataset. For evaluation purposes, single-word responses are generated. We assess the model’s accuracy and response quality (scored on a scale from 0 to 5) using GPT-4o-mini [33]. We also utilize the PerceptionTest [34] to assess perception capabilities, alongside VideoMME [8] and NeXTQA [46], which encompass various video domains and durations. VideoDetailCaption [31] assesses the model’s detailed understanding of video content, with scoring also conducted via GPT-4o-mini [33]. Additionally, we further evaluate the model on the MVbench [23] dataset, which includes 20 complex tasks requiring comprehensive video understanding beyond single-frame analysis. Each task contains 200 test samples in a multiple-choice VideoQA format. These samples require the model to choose the correct answer from multiple provided options. For MVBench, we conducted multiple experiments and computed the average results.
@@
Comparison Methods. We compare our method with two latest training-free visual token compression methods: LLaVA-PruMerge [39] leverages the sparsity of attention scores in CLIP [35] to identify key tokens; FastV [3] adopts the attention score between the predicted tokens and the visual tokens during the prefilling stage to identify key tokens. Notably, FastV offers two versions, with and without KV cache. In practice, we found that FastV without KV cache causes a sharp increase in inference time for video input, thereby we mainly compare to the version with KV cache. Of note, the above two methods rely on one-shot token pruning, while our approach introduces dynamic visual token pruning for the first time. We use the official codes [^1] https://github.com/pkunlp-icler/FastV, https://github.com/42Shawn/LLaVA-PruMerge of these methods for evaluation under identical hardware conditions.
@@
Implementation Details. We implement the proposed DyCoke on the LLaVA-OneVision-0.5B, LLaVA-OneVision-7B, and LLaVA-OneVision-72B models using NVIDIA RTX 4090 (24GB), A6000 (48GB), and A100 (80GB) GPUs, respectively. And we use the PyTorch framework. To set the pruning ratios for all methods, we use total calculated FLOPs to ensure fair comparison. The attention computation layer for FastV is set to layer 5. For video input, we follow the official requirements of the LLaVA-OneVision model, with a default of 32 video input frames and $N_{v}=196$, except for experiments with specific instructions. In the comparison experiments, $L$ is set to 3 and $P$ to 0.7, with the first-stage pruning rate $K$ serving as the primary experimental variable. For benchmarks such as PerceptionTest [34], VideoMME [8], NeXTQA, VideoDetailCaption [31], and ActivityNet-QA [50], we use the LMMs-Eval [53, 17] for evaluation, while MVBench [23] is evaluated using the official code.
@@
Method Total Latency $\downarrow$ GPU Mem. $\downarrow$ Accuracy $\uparrow$ Latency per Example $\downarrow$ LLaVA-OV-7B Full Tokens 1:19:27 27G 57.58 (±$(\pm 0.11)$ 1.19s (1.00$\times$) PruMerge 2:02:17 20G 52.59 (±$(\pm 0.08)$ 1.83s (0.65$\times$) FastV 1:03:50 24G 56.05 (±$(\pm 0.13)$ 0.96s (1.24$\times$) Ours ($K=0.5$) 59:02 21G 57.88 (±$(\pm 0.10)$ 0.88s (1.35$\times$) Ours ($K=0.7$) 57:13 19G 57.45 (±$(\pm 0.18)$ 0.85s (1.40$\times$) LLaVA-OV-0.5B Full Tokens 34:05 19G 47.09 (±$(\pm 0.09)$ 0.56s (1.00$\times$) PruMerge 1:35:08 8.5G 42.53 (±$(\pm 0.30)$ 1.42s (0.41$\times$) FastV 32:30 10G 46.14 (±$(\pm 0.03)$ 0.49s (1.14$\times$) Ours ($K=0.5$) 32:17 8.9G 46.76 (±$(\pm 0.13)$ 0.48s (1.16$\times$) Ours ($K=0.7$) 31:45 7.4G 47.09 (±$(\pm 0.11)$ 0.47s (1.19$\times$) (a) The number of video input frames is 16 Method Total Latency $\downarrow$ GPU Mem. $\downarrow$ Accuracy $\uparrow$ Latency per Example $\downarrow$ LLaVA-OV-7B Full Tokens 2:33:30 34G 58.81 (±$(\pm 0.14)$ 2.30s (1.00$\times$) PruMerge 3:48:59 28G 53.93 (±$(\pm 0.06)$ 3.43s (0.64$\times$) FastV 1:55:20 30G 58.36 (±$(\pm 0.09)$ 1.73s (1.32$\times$) Ours ($K=0.5$) 1:53:17 28G 59.09 (±$(\pm 0.23)$ 1.69s (1.36$\times$) Ours ($K=0.7$) 1:39:49 24G 59.56 (±$(\pm 0.19)$ 1.49s (1.54$\times$) LLaVA-OV-0.5B Full Tokens 1:19:06 21G 48.23 (±$(\pm 0.15)$ 1.18s (1.00$\times$) PruMerge 3:06:06 13G 42.55 (±$(\pm 0.18)$ 2.79s (0.42$\times$) FastV 1:13:25 15G 47.04 (±$(\pm 0.05)$ 1.10s (1.07$\times$) Ours ($K=0.5$) 1:10:32 12G 48.13 (±$(\pm 0.09)$ 1.06s (1.11$\times$) Ours ($K=0.7$) 1:02:15 10G 47.77 (±$(\pm 0.10)$ 0.93s (1.27$\times$) (b) The number of video input frames is 32